• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 03:24:29 on March31,2005

    Entry number 142850

    keyword=Current from 45->40 uA

    MCC called and asked if they could drop the beam current from 45 uA to 40 uA 
    due to vacuum problems.   We said this is fine.  The change occurred during run